Solve the following initial value problem:

\frac{dy}{dx}=x-\frac{1}{x^2} (x>0), y(1)=0
Answer #1

dy y = x - 12 22- dy = (x - 11/2) dx Jay= (pet) dx Sede af 2 de y = XP ya x² ~2+1 x -2+1 c is constant + c 2 2 = ๕// 주+ of c